OpenStudy (anonymous):

A bacterial culture contains 5000 bacteria, and the number of bacteria doubles each day. How many bacteria will be in the culture after a week? F(x)= p*a^x

OpenStudy (anonymous):

\[5000\times 2^7\] and a calculator

OpenStudy (anonymous):

i have to solve it step by step . so basically its like f(x)=5000*A^x f(x)=5000 *2^7 Is that right hun?

OpenStudy (anonymous):

yes darling

OpenStudy (anonymous):

 lol idk what goes in the boxes